OpenStudy (anonymous):

If f(x) = x^2 -1 and g(x) = 1/x +1 , the value of f(1/g(x)) is

OpenStudy (anonymous):

I am getting a different answer to the correct answer given

OpenStudy (anonymous):

@ganeshie8 @Kainui

OpenStudy (anonymous):

can anyone just tell me what answer are they getting

ganeshie8 (ganeshie8):

\(g(x) = \dfrac{1}{x} + 1\) ?

OpenStudy (anonymous):

I did all that, i am getting a different answer

OpenStudy (anonymous):

oh yes

ganeshie8 (ganeshie8):

\(g(x) = \dfrac{1}{x} + 1 = \dfrac{1+x}{x}\)

ganeshie8 (ganeshie8):

\(\large \implies \dfrac{1}{g(x)} = \dfrac{x}{1+x}\)

OpenStudy (anonymous):

yes got that

OpenStudy (anonymous):

There are options given wait

ganeshie8 (ganeshie8):

okie

OpenStudy (anonymous):